OpenStudy (anonymous):

Find the vertex of the parabola formed by the following quadratic function: y=-3x^2+12x-7 A)(-2,-43) B)(2,5) C)(-1,-22) D)(1,2)

OpenStudy (misty1212):

lol yeah dear why not?

OpenStudy (misty1212):

it is the same as the last one, use \[-\frac{b}{2a}\]

OpenStudy (misty1212):

in your example \(b=12,a=-3\) what do you get?

OpenStudy (anonymous):

idk I don't think im doing it right

OpenStudy (anonymous):

or the calc on my phone is doing something wrong

OpenStudy (misty1212):

lol use paper dear, it is easier

OpenStudy (misty1212):

\[-\frac{12}{2\times (-3)}\]

OpenStudy (misty1212):

two minus signs makes a plus, and 6 goes in to 12 2 times